A circular area with a radius of 7.70 cm lies in the x-y plane.

What is the magnitude of the magnetic flux through this circle due to a uniform magnetic field B = 0.212 T that points at an angle of 52.6 ∘ from the+z direction?

What is the magnitude of the magnetic flux through this circle due to a uniform magnetic field B = 0.212 T that points in the +y direction?
Answer #1

\Phi =BScos\alpha

\alpha=\measuredangle \left ( \vec{B}, \vec{S} \right )

Vector S is considered by its normal to the area.

In the first case: \Phi =BScos\left ( \pi -\alpha \right )=-BScos\alpha =-B\pi r^{2}cos\alpha\approx 24\: Wb

In the second case: the angle between vector B and the normal to the surface is 90 deg. cos90^{\circ}=0

Thus, \Phi =0\: Wb in this case.